Solve the following motion problems. Set up an equation and solve.
A boy rode his bicycle into the country at the rate of 15 miles per hour and returned over the same route at 12 miles per hour. If he was bicycling for 4.5 hours, how far into the country did he ride?

Answer by checkley77(12844)

D=RT
12T=15(4.5-T)
12T=67.5-15T
12T+15T=67.5
27T=67.5
T=67.5/27
T=2.5 HOURS FOR THE 12 MPH. TRIP.
12*2.5=30 MILES IS THE DISTANCE.
15(4.5-2.5)=15*2=30 MILES DITTO.